# Feedback Descent: Open-Ended Text Optimization via Pairwise Comparison

### v1

@@ 2.3 Analogy to Gradient Descent
We emphasize that this gradient analogy serves only as an intuition: gradients indicate directions in continuous parameter space, while rationales indicate directions in semantic space, without guarantees of smoothness or convergence. Our experiments in Section˜5 demonstrate that feedback-guided optimization indeed produces consistent improvements across tasks, validating that such heuristic directional cues are sufficient to drive open-ended text optimization.
3 Directional Information is Critical When Effective Dimensionality is High
In this section, we give intuition for why and when Feedback Descent is particularly effective, using a minimal idealized model of directional information based on comparisons with feedback. We provide simplified statements here, and show full proofs in Appendix˜A.
Pairwise preference alone is a binary signal: it tells us which artifact is better but not how to improve. Adding a short rationale (“legs disconnected from body; increase stroke width”) turns each comparison into a directional cue in semantic space. Iterating this loop accumulates such cues, allowing a mutator LLM to make small, targeted edits that are more likely than chance to increase quality. Even when individual rationales are imperfect, the aggregate message across failures (what kept losing and why) continually refines the direction of improvement.
As an idealized model, map each text $x$ to a latent feature $z=\phi(x)\in\mathbb{R}^{d}$ with utility $r(z)=R(x)$, where $d$ denotes the effective dimensionality—the number of independent directions along which artifacts can vary meaningfully. Let $z^{\star}=\arg\max_{z}r(z)$ be the optimum. For problems of interest, $d$ may be in the hundreds or thousands, reflecting the many degrees of freedom in text, molecule, or image design spaces.
A rationale $r_{t}$ is useful if its induced edit defines a direction $v_{t}$ positively correlated with the gradient $\nabla r(z_{t})$. Under this model, feedback carries gradient information and systematically reduces error, yielding linear convergence in expectation, even with imperfect rationales.
Proposition 1 (Directional updates achieve dimension-free convergence, informal).
Suppose we perform $N$ rationale-guided updates, where each update direction $v_{t}$ is on average positively aligned with $\nabla r(z_{t})$. Then the expected error after $N$ evaluations satisfies
$\mathbb{E}[\,r(z^{\star})-r(z_{N})\,]\;\leq\;(1-\alpha)^{N}\,[r(z^{\star})-r(z_{0})],$
for some constant $\alpha>0$ depending only on the alignment and variability of the feedback. Crucially, this rate does not depend on the dimension $d$.
While rationale-guided updates exploit directional information to achieve convergence rates that do not depend on the ambient dimension, methods that rely only on function evaluations or binary preferences suffer severe dimension-dependent slowdowns:
Proposition 2 (Inefficiency of zeroth-order search, informal).
Consider zeroth-order methods with $N$ evaluations in $\mathbb{R}^{d}$. For a strongly concave quadratic, to reach $\epsilon$-accuracy one needs at least $N\;\geq\;(c/\epsilon)^{d/2}$ and best-of-$N$ random sampling achieves at best
$\mathbb{E}[\,r(z^{\star})-r(z_{N})\,]=\Omega(N^{-2/d}).$
Thus, while rationale-guided updates drive $r(z_{N})$ toward $r(z^{\star})$ at a dimension-free linear rate, zeroth-order search suffers exponential slowdowns as $d$ grows.
We emphasize that these propositions are highly idealized, serving only to illustrate a general principle: in high-dimensional search, even weak directional cues can accumulate into steady improvement, whereas unguided search scales exponentially worse. Thinking in terms of a continuous latent space is a pragmatic abstraction; modern ML operates on continuous weights and embeddings that approximate local structure even in discrete domains. By the same logic, many forms of human progress, including writing, coding, and science, often arise from successive refinements within an implicit semantic space shared by experts. The following statements are for motivational clarity, not literal models of the systems in Section˜5.
The propositions rest on well-known results in optimization and combinatorics. Our contribution is applying this intuition to open-ended text optimization, where it highlights effective dimensionality as the key bottleneck. To our best knowledge, the closest prior work that adopts this framing is the empirical taxonomy and optimizer design of nie2024importance; building on their study, we identify effective dimensionality as the key limiting factor, showing that in open-ended settings where artifacts admit many possible alterations, even weak but aligned textual rationales provide the directional information needed for linear convergence, while zeroth-order methods degrade exponentially.
Implications. First, even imperfect cues accelerate optimization when aggregated, since their consistent positive correlation with the gradient accumulates into reliable directional information. Second, in problems with high effective dimensionality, where exploration can proceed in unbounded directions, pure zeroth-order search becomes extremely inefficient. In this regime, obtaining directed feedback on the most salient aspects, whether from experimental results, an expert’s feedback, or exploiting a generator-verifier gap, is essential for sustained progress. This is especially important for open-ended problems where the optimum is not known in advance.
Directional Information is Advantageous in High Dimensions Unlike zeroth-order methods, which scale poorly with dimension, Feedback Descent remains effective when the effective dimensionality is high because feedback provides directional cues that guide optimization.
@@ 4 Related Work
Gradient-Free Optimization. Zeroth-order optimization methods [chen2019zoadamm] operate without access to gradients, typically achieving convergence rates that scale poorly with dimension. While the black-box nature of modern LLMs has spurred interest in applying gradient-free approaches [han2023match, sun2024bbtv2, jin2024derivative, brahmachary2024evolution], these methods face fundamental challenges in high-dimensional spaces. Our approach explores whether textual rationales can provide useful directional information for optimization, similar to how nie2024importance shows that LLMs can be effective optimizers when provided with directional feedback from historical traces.
Optimizing Compound AI Systems. Compound AI systems, i.e., modular architectures involving multiple LLM invocations and complex control flow, such as agents or scaffolding techniques [yao2022react], present unique optimization challenges due to their modularity. Several approaches have emerged to tackle this complexity, including optimization for searching and bootstrapping few-shot in-context examples [khattab2022dspy, khattab2024dspy, opsahl2024optimizing] and reflective prompt evolution [agrawal2025gepa]. The closest prior work to Feedback Descent is TextGrad [yuksekgonul2024textgrad], which proposes PyTorch-like framework for automatic differentiation via text, backpropagating textual gradients through computation graphs. A core difference is that TextGrad optimizes “pointwise”, i.e., it proposes a new instance of the artifact based only on the latest one. In contrast, Feedback Descent keeps an explicit trajectory-level buffer of comparative feedback. As we will see in Section˜5, Feedback Descent is much more scalable at long-horizon optimization compared to TextGrad.
@@
We build on the growing consensus that natural language is a particularly powerful medium for inference-time improvement. Natural language traces enable models to reason effectively in complex environments [lampinen2022tell, wei2022chain], and language models can reliably map textual instructions to concrete modifications [chen2021evaluating, austin2021program, saunders2022self, scheurer2023training]. However, existing methods often rely on random sampling of self-generated critiques, which may be noisy or fail to capture external preferences. In contrast, we leverage external rationales as directional information, enabling guided search in the semantic space.
@@ 5 Experiments
We evaluate Feedback Descent across three diverse domains—visual design, prompt optimization, and molecule discovery—to demonstrate its generality and effectiveness. Our evaluation spans diverse representations and evaluation modalities: (1) Representation diversity: SVG (spatial/geometric), prompts (instructional text), molecules (chemical structures). (2) Evaluation modality: SVG uses vision-language model comparison, prompt optimization uses dataset-specific accuracy metrics, molecules use computational docking scores. Together, these answer: Can a single optimization framework, with no domain-specific engineering, match or exceed specialized methods purely through structured feedback?
@@ 5.1 Experimental Domains
Setup Subject Ink Wash Minimalist Realism Retro Arcade Stained Glass
From Scratch City Skyline 100.0% 100.0% 100.0% 100.0% 100.0%
Mushroom 100.0% 100.0% 83.3% 100.0% 100.0%
Robot 100.0% 100.0% 100.0% 100.0% 100.0%
Unicorn 100.0% 100.0% 100.0% 100.0% 100.0%
Judge-Aware City Skyline 50.0% 100.0% 100.0% 100.0% 88.9%
Mushroom 100.0% 100.0% 100.0% 100.0% 94.1%
Robot 100.0% 100.0% 100.0% 93.8% 100.0%
Unicorn 89.5% 100.0% 100.0% 93.8% 80.0%
Table 1: Win rates after five iterations comparing Feedback Descent against direct prompting under two conditions: from Scratch and Informed of the judge rubric. Values above 50% indicate that Feedback Descent outperforms the baseline; denotes the only failure case where direct prompting wins. Iterative feedback consistently improves SVG designs over direct prompting.
@@ 5.2 SVG Optimization
Table 2: Prompt optimization results across multiple benchmarks. GEPA consistently outperforms baseline methods across different model architectures.
Results. Table˜1 shows the win rates after 5 iterations. For both GPT-4o-mini and GPT-5-mini, Feedback Descent reliably improves outputs over the initial population. Furthermore, qualitative examples in Figure˜3 demonstrate that the procedure consistently produces unicorns whose visual style diverges across judges, aligning with aesthetic criteria such as geometry, minimalism, or retro arcade motifs.
@@ 5.3 Prompt Optimization
Pairwise Comparisons Suffice for Competitive Prompt Optimization Feedback Descent remains competitive with GEPA, a state-of-the-art prompt optimization method, despite employing a simpler, domain-agnostic approach that focuses on summarizing pairwise performance differences.
Method ADRB1 PGR PPARA PPARG CDK2 F2
DOCKSTRING (N=260155)
Top $50\%$ 5.305 3.478 4.549 4.210 4.385 4.168
Top $90\%$ 8.785 7.878 7.987 7.658 7.733 7.477
Top $99\%$ 9.620 8.703 8.718 8.449 8.453 8.139
Top $99.9\%$ 10.209 9.260 9.230 9.012 8.979 8.722
Top $99.99\%$ 10.742 9.723 9.821 9.518 9.509 9.252
Best Molecule 11.330 9.742 9.907 9.529 9.534 9.311
GP-BO [tripp2021a] 10.552 9.307 9.680 9.485 9.067 8.686
Graph MCTS [jensen2019graph] 8.883 7.819 7.363 7.134 7.777 6.310
Graph GA [jensen2019graph] 9.145 8.670 8.598 8.327 8.288 8.102
SMILES GA [brown2019guacamol] 9.334 8.335 9.052 8.560 8.268 7.984
REINVENT [olivecrona2017molecular] 9.018 8.267 8.430 8.347 8.226 8.139
TextGrad [yuksekgonul2024textgrad] 8.531 8.057 7.953 7.256 8.174 7.357
Feedback Descent 10.623 9.615 9.919 10.187 9.803 9.300
Table 3: Comparison of molecule optimization methods on six protein targets. Fragment-based algorithms (denoted by $\dagger$) operate directly on molecular graphs, giving them structural priors unavailable to purely text-based methods. For each target, the top generative result is in bold, and any population in the DOCKSTRING that exceeds the best generative result is underlined. Feedback Descent rivals or surpasses specialized molecular optimizers across all six targets.
@@ 5.4 Molecule Optimization (DOCKSTRING)
Results. Table˜3 summarizes optimization outcomes across six protein targets. Feedback Descent outperforms all baselines and achieves the strongest scores on all targets. On multiple proteins, it matches or exceeds the 99.9th and even 99.99th percentiles of the DOCKSTRING database, including surpassing the best molecule present in the database itself ($N=260155$). Notably, TextGrad consistently underperforms Feedback Descent across all targets; while the TextGrad paper evaluated on similar DOCKSTRING tasks for only $\sim$10 iterations, we run both methods for 1000 steps and find that TextGrad’s pointwise conditioning does not scale well to high iteration budgets, confirming that accumulated feedback history is essential for sustained improvement. These findings show that Feedback Descent, a purely text-based method, can rival or outperform specialized graph-based algorithms, despite lacking handcrafted structural priors.
Figure˜5 shows optimization trajectories for PPARG. Feedback Descent achieves competitive trajectories relative to specialized methods, reaching high-scoring regions of chemical space with comparable or fewer oracle calls. This pattern holds across targets, suggesting that the method generalizes rather than relying on idiosyncrasies of a single protein system.
Analysis of discovered molecules. Figure˜6 illustrates the Pareto frontier between docking affinity (Vina score) and drug-likeness (QED) for PPARG. Feedback Descent recovers molecules that sit on or above the DOCKSTRING frontier, indicating that improvements in affinity are achieved without compromising drug-likeness. See Figure˜7 in the appendix for the full set of Pareto frontiers across all targets. These results show that feedback-guided search yields candidates that are not only potent but also balanced along multiple drug-relevant dimensions.
We also examine novelty by plotting Tversky similarity (CFP4 fingerprints) to approved DrugBank molecules against docking scores in Figure˜4. Across all targets, the correlations are weak or negative (Spearman $\rho$ between 0.39-0.39 and $0.40$), indicating that high-scoring candidates identified by Feedback Descent do not simply recycle functional groups from existing drugs, but instead explore novel regions of chemical space. For CDK2, no comparison is shown, as the target lacks any fully approved drugs in DrugBank with orthosteric binding as part of their mechanism of action, and thus does not satisfy our inclusion criteria.

### v2

@@ 2.3 Analogy to Gradient Descent
Why directional information helps. Zeroth-order methods that rely only on function evaluations or binary preferences suffer severe dimension-dependent slowdowns: convergence rates degrade exponentially as the search space grows [nemirovski1983problem, nesterov2017random]. In contrast, first-order methods exploit gradient information to achieve dimension-free convergence under standard assumptions. Textual feedback provides an approximation to such directional information. Even when individual rationales are imperfect, their aggregate message across failures continually refines the direction of improvement. We formalize this intuition in Appendix˜A, showing that under idealized assumptions, rationale-guided updates can achieve linear convergence rates independent of effective dimensionality, while zeroth-order baselines scale exponentially worse. These results provide motivation rather than rigorous guarantees for the discrete text domains we study empirically. In Section˜4, we show that Feedback Descent indeed produces consistent improvements across tasks, validating that such heuristic directional cues are sufficient to drive open-ended text optimization.
@@ 3 Related Work
Evolutionary Algorithms and Gradient-Free Optimization. Feedback Descent can be viewed as an evolutionary algorithm [golberg1989genetic, holland1992adaptation], in which candidates are iteratively mutated and accepted based on fitness. While the black-box nature of modern LLMs has spurred interest in applying gradient-free approaches [guo2023match, sun2024bbtv2, chen2024derivative, lange2024evolution], these methods face fundamental challenges in high-dimensional spaces. More broadly, zeroth-order methods [chen2019zoadamm] face convergence rates that scale poorly with dimension, which is consistent with our experimental results comparing with reinforcement learning methods in Section˜4. To our best knowledge, wang2024efficient was the first prior work to demonstrate that LLMs can be effective molecule optimizers inside evolutionary algorithm loops. Feedback Descent explores whether textual rationales can provide useful directional information for optimization, similar to how nie2024importance shows that LLMs can be effective optimizers when provided with directional feedback from historical traces. Our contribution is in operationalizing an effective directed mutation operator via accumulated textual feedback.
Optimizing Compound AI Systems. Compound AI systems, i.e., modular architectures involving multiple LLM invocations and complex control flow, such as agents or scaffolding techniques [yao2022react], present unique optimization challenges due to their modularity. Several approaches have emerged to tackle this complexity, including optimization for searching and bootstrapping few-shot in-context examples [khattab2022dspy, khattab2024dspy, opsahl2024optimizing] and reflective prompt evolution [agrawal2025gepa]. The closest prior work to Feedback Descent is TextGrad [yuksekgonul2024textgrad], which proposes PyTorch-like framework for automatic differentiation via text, backpropagating textual gradients through computation graphs. A core difference is that TextGrad optimizes “pointwise”, i.e., it proposes a new instance of the artifact based only on the latest one. In contrast, Feedback Descent keeps an explicit trajectory-level buffer of comparative feedback. As we will see in Section˜4, Feedback Descent is much more scalable at long-horizon optimization compared to TextGrad.
@@
We build on the growing consensus that natural language is a particularly powerful medium for inference-time improvement. Natural language traces enable models to reason effectively in complex environments [lampinen2022tell, wei2022chain], and language models can reliably map textual instructions to concrete modifications [chen2021evaluating, austin2021program, saunders2022self, scheurer2023training, feng2024natural]. However, existing methods often rely on random sampling of self-generated critiques, which may be noisy or fail to capture external preferences. In contrast, we leverage external rationales as directional information, enabling guided search in the semantic space.
@@ 4 Experiments
We evaluate Feedback Descent across three diverse domains—visual design, prompt optimization, and molecule discovery—to demonstrate its generality and effectiveness. Our evaluation spans diverse representations and evaluation modalities: (1) Representation diversity: SVG (spatial/geometric), prompts (instructional text), molecules (chemical structures). (2) Evaluation modality: SVG uses vision-language model comparison, prompt optimization uses dataset-specific accuracy metrics, molecules use computational docking scores. Together, our experiments answer the question: Can a single optimization framework, with no domain-specific engineering, match or exceed specialized methods purely through structured feedback?
@@ 4.1 Experimental Domains
Setup Subject Ink Wash Minimalist Realistic Arcade Stain Glass
From Scratch City Skyline 100.0100.0% 100.0100.0% 100.0100.0% 100.0100.0% 100.0100.0%
Mushroom 100.0100.0% 100.0100.0% 83.383.3% 100.0100.0% 100.0100.0%
Robot 100.0100.0% 100.0100.0% 100.0100.0% 100.0100.0% 100.0100.0%
Unicorn 100.0100.0% 100.0100.0% 100.0100.0% 100.0100.0% 100.0100.0%
Judge-Aware City Skyline 50.050.0% 100.0100.0% 100.0100.0% 100.0100.0% 88.988.9%
Mushroom 100.0100.0% 100.0100.0% 100.0100.0% 100.0100.0% 94.194.1%
Robot 100.0100.0% 100.0100.0% 100.0100.0% 93.893.8% 100.0100.0%
Unicorn 89.589.5% 100.0100.0% 100.0100.0% 93.893.8% 80.080.0%
Table 1: Win rates after five iterations comparing Feedback Descent against direct prompting under two conditions: From Scratch and Aware of the judge rubric. Feedback Descent matches or outperforms the baseline on all combinations tested (i.e. $\geq 50\%$). Iterative feedback consistently improves SVG designs over direct prompting.
@@ 4.2 SVG Optimization
Table 2: Prompt optimization results across multiple benchmarks. Feedback Descent consistently outperforms or is competitive with state-of-the-art methods.
Results. Table˜1 shows the win rates after 5 iterations. For both GPT-4o-mini and GPT-5-mini, Feedback Descent reliably improves outputs over the initial population. Furthermore, qualitative examples of an optimization trajectory in Figure˜2 and each judge-object pair in Figure˜3 demonstrate that the procedure consistently produces unicorns whose visual style diverges across judges, aligning with aesthetic criteria such as geometry, minimalism, or retro arcade motifs.
@@ 4.3 Prompt Optimization
Pairwise Comparisons Suffice for Competitive Prompt Optimization Feedback Descent is competitive with GEPA, a state-of-the-art prompt optimization method, despite employing a simpler, domain-agnostic approach.
Method ADRB1 PGR PPARA PPARG CDK2 F2 Avg
DOCKSTRING (N=260155) Top $50\%$ $5.305$ $3.478$ $4.549$ $4.210$ $4.385$ $4.168$ $4.349$
Top $90\%$ $8.785$ $7.878$ $7.987$ $7.658$ $7.733$ $7.477$ $7.920$
Top $99\%$ $9.620$ $8.703$ $8.718$ $8.449$ $8.453$ $8.139$ $8.680$
Top $99.9\%$ $10.209$ $9.260$ $9.230$ $9.012$ $8.979$ $8.722$ $9.235$
Top $99.99\%$ 10.742 9.723 $9.821$ $9.518$ $9.509$ $9.252$ $9.761$
Best Molecule 11.330 9.742 $9.907$ $9.529$ $9.534$ 9.311 $9.892$
Graph MCTS [jensen2019graph] $8.883$ $7.819$ $7.363$ $7.134$ $7.777$ $6.310$ $7.548$
SMILES GA [brown2019guacamol] $9.334$ $8.335$ $9.052$ $8.560$ $8.268$ $7.984$ $8.589$
REINVENT [olivecrona2017molecular] $9.867$ $8.604$ $8.735$ $9.054$ $8.695$ $8.441$ $8.899$
Graph GA [jensen2019graph] $10.249$ $8.793$ $9.211$ $8.769$ $8.652$ $8.900$ $9.096$
GP-BO [tripp2021a] $10.552$ $9.307$ $9.680$ $9.485$ $9.067$ $8.686$ $9.463$
TextGrad [yuksekgonul2024textgrad] $8.531$ $8.057$ $7.953$ $7.256$ $8.174$ $7.357$ $7.888$
Feedback Descent (ours) $10.623$ $9.615$ $9.919$ $10.187$ $9.803$ $9.300$ $9.908$
    w/ No Feedback $6.190$ $8.619$ $8.230$ $8.633$ $8.300$ $8.793$ $8.127$
    w/ Random Feedback $6.604$ $8.385$ $8.276$ $6.780$ $8.793$ $7.993$ $7.805$
    w/ Binary Only $5.863$ $8.779$ $8.507$ $7.998$ $9.439$ $8.420$ $8.168$
Table 3: Results for molecule optimization on six protein targets. Full results with standard deviations are in Table˜6. For each target, the top generative result is in bold, and any population in the DOCKSTRING database that exceeds the best generative result is underlined. Feedback Descent rivals or surpasses specialized molecular optimizers across all six targets.
@@ 4.4 Molecule Optimization (DOCKSTRING)
Noise ADRB1 PGR PPARG
None $10.62$ $9.62$ $10.19$
25% $9.28$ $9.14$ $8.16$
50% $10.21$ $8.92$ $8.75$
100% $6.60$ $8.39$ $6.78$
Table 4: DOCKSTRING scores for ADRB1, PGR, and PPARG with Feedback Descent at varying feedback noise levels. Performance degrades gracefully with increasing noise.
@@
Main Results.
Target Win Rate
ADRB1 76%
CDK2 86%
F2 79%
PPARG 83%
Avg 81%
Table 5: True feedback vs. scrambled feedback win rate across 400 comparisons.
Table˜3 summarizes optimization outcomes across six protein targets. Feedback Descent outperforms all baselines and achieves the strongest scores on all targets. On multiple proteins, it matches or exceeds the 99.9th and even 99.99th percentiles of the DOCKSTRING database, including surpassing the best molecule present in the database itself ($N=260155$). Notably, TextGrad consistently underperforms Feedback Descent across all targets; while the TextGrad paper evaluated on similar DOCKSTRING tasks for only $\sim$10 iterations, we run both methods for 1000 steps and find that TextGrad’s pointwise conditioning does not scale well to high iteration budgets, confirming that accumulated feedback history is essential for sustained improvement. These findings show that Feedback Descent, a purely text-based method, can rival or outperform specialized graph-based algorithms, despite lacking handcrafted structural priors.
Figure˜4 shows optimization trajectories for PPARG. Feedback Descent achieves competitive trajectories relative to specialized methods, reaching high-scoring regions of chemical space with comparable or fewer oracle calls. This pattern holds across targets, suggesting that the method generalizes rather than relying on idiosyncrasies of a single protein system.
Analysis of discovered molecules. Figure˜5 illustrates the Pareto frontier between docking affinity (Vina score) and drug-likeness (QED) for PPARG. Feedback Descent recovers molecules that sit on or above the DOCKSTRING frontier, indicating that improvements in affinity are achieved without compromising drug-likeness. See Figure˜7 in the appendix for the full set of Pareto frontiers across all targets. These results show that feedback-guided search yields candidates that are not only potent but also balanced along multiple drug-relevant dimensions.
We also examine novelty by plotting Tversky similarity (CFP4 fingerprints) to approved DrugBank molecules against docking scores in Figure˜6. Across all targets, the correlations are weak or negative (Spearman $\rho$ between 0.39-0.39 and $0.40$), indicating that high-scoring candidates identified by Feedback Descent do not simply recycle functional groups from existing drugs, but instead explore novel regions of chemical space. For CDK2, no comparison is shown, as the target lacks any fully approved drugs in DrugBank with orthosteric binding as part of their mechanism of action, and thus does not satisfy our inclusion criteria.
@@
Feedback quality and type. In addition to the main points of comparison, we show three ablations on the quality of the feedback in Table˜3. We evaluate (1) No Feedback, i.e., parallel best-of-N sampling with no feedback; (2) Random Feedback, which uses the same iterative algorithm but shuffles rationales between molecule pairs; and (3) Binary Only, which provides only the binary signal of whether the newest molecule is better than the previous candidate. Random feedback underperforms even best-of-N, confirming that the method relies on feedback content rather than just the iterative structure. The gap between Binary Only and Feedback Descent (1.74 docking score units) demonstrates that textual rationales provide substantial value beyond the binary preference signal. Table˜4 further tests robustness to corrupted feedback by randomly shuffling rationales at varying noise levels. The method degrades gracefully: 50% noise still achieves scores above the 99.9th percentile of DOCKSTRING for ADRB1 and PGR.
Feedback alignment. Finally, we test whether generated molecules actually reflect the feedback shown during optimization in Table˜5. We present an LLM judge with a newly generated molecule and ask it to determine whether the molecule more closely follows the true feedback or a scrambled control. We randomly flip the order shown to the judge to account for order bias. Across 400 comparisons, true feedback wins 81% of head-to-head comparisons ($p<10^{-10}$, binomial test), demonstrating that the generator reliably reads and incorporates structured property feedback.